### v2

# for Human Motion Generation

Abstract
Human motion generation plays a vital role in applications such as digital humans and humanoid robot control. However, most existing approaches disregard physics constraints, leading to the frequent production of physically implausible motions with pronounced artifacts such as floating and foot sliding. In this paper, we propose Morph, a Motion-free physics optimization framework, comprising a Motion Generator and a Motion Physics Refinement module, for enhancing physical plausibility without relying on costly real-world motion data. Specifically, the Motion Generator is responsible for providing large-scale synthetic motion data, while the Motion Physics Refinement Module utilizes these synthetic data to train a motion imitator within a physics simulator, enforcing physical constraints to project the noisy motions into a physically-plausible space. These physically refined motions, in turn, are used to fine-tune the Motion Generator, further enhancing its capability. Experiments on both text-to-motion and music-to-dance generation tasks demonstrate that our framework achieves state-of-the-art motion generation quality while improving physical plausibility drastically.
1 Introduction
Accurate human motion generation is pivotal in various applications, including robotics, video gaming and virtual reality [10, 36, 45, 44, 1]. Recent advances in AI have paved the way for novel approaches to motion generation, enabling various control conditions including textual descriptions and music pieces. Mainstream approaches can be divided into three categories: conditional diffusion models [32, 42, 39, 33, 16], conditional autoregressive models [23, 13, 46, 47, 18] and generative masked modeling [7, 24]. These motion generation models have made significant progress in capturing the complex multimodal distribution of human motions, greatly improving motion generation quality.
However, most existing motion generation approaches overlook a fundamental aspect of human motion: the laws of physical. Although these generation models excel in capturing statistical distribution of human motions, they lack explicit mechanisms to enforce physical constrains. As a result, the generated motions frequently exhibit pronounced artifacts such as ground penetration, leaning backward, interpenetration, foot sliding, floating and unnatural rotation, as shown in Fig. 1. Given human sensitivity to even slight physical inconsistencies, these physically implausible motions hinder many real-world applications such as animation and virtual reality [39, 27, 12].
Figure 1: Examples of physical inconsistencies in generations.
Recently, a few studies [39, 11] have attempted to enhance the physical plausibility of motion generation. For example, PhysDiff [39] and Reindiffuse [11] incorporate physical constraints into the denosing diffusion process. However, these approaches face several limitations. First, these works [39, 11] customize physics optimization operation for a specific type of generation model, i.e., text-conditional diffusion model. This customization restricts their transferability to a broader range of generation models (such as autoregressive-based models) and motion generation tasks (such as music-to-dance). Second, these works apply physics and diffusion iteratively, embedding physics optimization into multiple steps of diffusion process. The frequent execution of physics optimization significantly increases computational costs during inference. Last but most importantly, these methods require large-scale, high-quality realistic motion data to train the physics optimization mechanism. However, collecting such realistic motion data is challenging. A natural question arises: Is it possible to learn an efficient, model-agnostic physical optimizer without relying on real motion data?
To achieve this, we propose a Motion-free physics optimization framework, namely Morph. As shown in Fig. 2, Morph consists of two main modules: a Motion Generator (MG) that can be any existing motion generator, and a Motion Physic Refinement (MPR) module for enhancing physical plausibility. Morph employs a two-stage optimization process. In the first stage, using large-scale synthetic noisy motion data produced by the motion generator, MPR module is optimized to project input motions into a physically-plausible space. Specifically, the MPR module designs a motion imitator that controls a character agent to mimic the given noisy motions within a physics simulator. The simulator enforces multiple physical constraints, effectively reducing artifacts such as floating and foot sliding. To further ensure the naturalness of simulated motions, MPR module introduces a motion discriminator to align the distribution of physics-refined motions with that of input motions. Feedback signals from both physics simulator and motion discriminator guide the optimization of the motion imitator via reinforcement learning. In the second stage, we observe that due to the high cost of collecting 3D motion data, existing generators typically rely on limited motion data for training, which restricts their generative capabilities. In this stage, by leveraging the trained MPR module, a large-scale set of physically plausible, high-quality motion data can be created to finetune the generator, further enhancing its ability to generate realistic motion. During inference, the finetuned motion generator and MPR module work in tandem to generate physics-plausible and high-quality motions.
Through above process, Morph effectively decouples physic optimization from generation model, requiring only a single step of physical refinement and without relying on real motion data. This results in an economical, efficient, and model-agnostic physical optimizer. We evaluate our Morph framework on two motion generation tasks: text-to-motion and music-to-dance generation. Since Morph is agnostic to specific instantiation of generation models, we test it with three types of generation models, i.e., diffusion-based, autoregressive-based and generative mask modeling. Extensive experiments demonstrate that Morph achieves significant improvements in physical error metrics, while also achieving competitive generation metrics across different generators and tasks, despite not being trained on real motion data.
Figure 2: An overview of the Morph framework. Morph comprises a Motion Generator and a Motion Physics Refinement module. Morph employs a two-stage training process: Motion Physics Refinement module training and Motion Generator fine-tuning. And a Imitation Selection Operation is employed to ensure the motion quality after physics refinement.
2 Related Work
Human Motion Generation.  Motion generation is a long-history task that can be conditioned on various signals, such as text description, music and action [48, 7, 40, 8, 29, 33, 37, 2, 4, 30, 43]. Our work specifically focuses on text-to-motion [8, 40, 41, 32, 7, 24] and music-to-dance generation [33, 29, 16, 20]. Mainstream approaches can be roughly divided into three categories: diffusion-based methods [32, 3, 42, 39, 33, 16], autoregressive models [40, 17, 13, 46, 47, 18, 29] and generative masked modeling [7, 24]. For instance, MDM [32] uses a transformer-based diffusion model with conditional text representations extracted from CLIP [26]. T2M-GPT [40] designs a conditional autoregressive transformer model based on VQ-VAE [34] and GPT [25, 35]. MoMask [7] employs residual vector quantization and generative masked transformers to iteratively generate motions. For music-to-dance generation, Bailando [29] predicts discrete token sequences conditioned on music and uses an autoregressive transformer to regenerate the dance sequence. However, existing motion generation models often produce physically implausible motions, as they overlook physical laws during training. In contrast, our proposed MPR module effectively enforces physical constraints, significantly enhancing the physically plausibility of generated motions.
Physically Plausible Motion Generation. Physical plausibility refers to the degree to which generated motions adhere to physical rules, such as foot sliding, foot-ground contact, and body leaning. Generating physically-plausible motions is crucial for many real-world applications, such as animation and virtual reality. Recently, a few works have attempted to address this challenge [39, 11, 33]. For example, [39, 11] integrate physical constraints into the diffusion process, iteratively applying physics and diffusion to maintain alignment with motion data distribution while enhancing physical realism. EDGE [33] introduces auxiliary losses to align specific aspects of physical realism, specifically for generating more physically plausible dances. However, these methods tailor physic optimization to diffusion models and require real motion data, limiting their flexibility and scalability. In contrast, we propose a model-agnostic, motion-free physics optimization framework that decouples physics optimization from the generation model and learns to enhance physical plausibility without real motion data, offering a more economical and versatile solution.
3 Method
To enhance physical plausibility of motion generation without relying on real motion data, we propose a motion-free physics optimization framework, namely Morph. As illustrated in Fig. 2, Morph consists of two key modules: a Motion Generator (MG) that could be any existing pre-trained motion generation model, and a Motion Physics Refinement module (MPR) independent of Motion Generator. Morph employs a two-stage training process. In the first stage, the Motion Generator produces large-scale noisy motion data (Sec. 3.1), which is subsequently used to train the Physics Refinement module to map input motions into a physically plausible space (Sec. 3.2). In the second stage, these physics-refined motions, in turn, are used to fine-tune Motion Generator, further enhancing its capabilities (Sec 3.3). Through this two-stage training process, Morph drastically reduces physical errors while improving the overall quality of generated motions.
3.1 Noisy Motion Data Generation
Different from  [39, 11] that rely on real motion data, our goal is to develop a motion-free physical optimization framework. To achieve this, we utilize an existing pretrained motion generation model to produce large-scale, noisy motion data for training the physics refinement module. As shown in Fig. 2, given control signals $c$ (e.g., text or music), the pre-trained motion generator is employed to generate a motion sequence $\bm{\tilde{x}}^{1:L}$, as follows:
$\bm{\tilde{x}}^{1:L}=f_{\xi}\left(c\right),\text{where},\bm{\tilde{x}}^{1:L}=\left\{\bm{\tilde{x}}^{l}=\left[\bm{\theta}^{l},\bm{p}^{l}\right]\right\}_{l=1}^{L},$ (1)
where $f_{\xi}$ represents the motion generator with parameters $\xi$, $\bm{\tilde{x}}^{l}$ denotes the $l^{th}$ pose of the synthetic motion sequence, represented by the joint rotations $\bm{\theta}$ and positions $\bm{p}$. Notably, our framework, Morph, is agnostic to the specific instantiation of $f_{\xi}$, allowing it to be generally compatible with various pretrained motion generators.
3.2 Physics-Based Motion Refinement
Since most existing motion generation models lack explicit physical constraints, they often produce motions with noticeable artifacts, as shown in Fig. 1. To address this, the Motion Physics Refinement (MPR) module is tasked with projecting the generated motion $\bm{\tilde{x}}^{1:L}$, which disregards the laws of physics, into a physically-plausible motion $\bm{\hat{x}}^{1:L}$. As shown in Fig. 2, the MPR module consists of three components: a motion imitator, a physics simulator and a motion discriminator. Specifically, the motion imitator controls a simulated character to mimic the input motion $\bm{\tilde{x}}^{1:L}$ within the physics simulator. The resulting motion $\bm{\hat{x}}^{1:L}$ from the physics simulator is considered physically plausible, as it adheres to the laws of physics. Additionally, the motion discriminator seeks to align the distribution of refined motions $\bm{\hat{x}}^{1:L}$ with that of input motions $\bm{\tilde{x}}^{1:L}$, further ensuring the naturalness and realism of the physics-refined motions.
Motion Imitator Learning.
Motion Imitator Learning can be formulated as a Markov decision process, represented by the tuple $\left(\mathcal{S},\mathcal{A},\mathcal{T},\mathcal{R}\right)$ of states, actions, transition dynamics and reward function. Formally, the motion imitator is described by a policy $\pi\left(\bm{a}^{l}|\bm{s}^{l}\right)$, which specifies the probability distribution of selecting an action $\bm{a}^{l}\in\mathcal{A}$ given the current state $\bm{s}^{l}\in\mathcal{S}$. The physic simulator, in turn, defines the transition dynamics $\mathcal{T}\left(\bm{s}^{l+1}|\bm{s}^{l},\bm{a}^{l}\right)$, which determines the next state $\bm{s}^{l+1}$ based on current state $\bm{s}^{l}$ and action $\bm{a}^{l}$. Specifically, starting from an initial state $\bm{s}^{1}$, a character agent acts in the physic simulator according to policy $\pi\left(\bm{a}^{l}|\bm{s}^{l}\right)$, iteratively sampling action $\bm{a}^{l}$. Then, the physic simulator, governed by the transition dynamics $\mathcal{T}\left(\bm{s}^{l+1}|\bm{s}^{l},\bm{a}^{l}\right)$, generates the next state $\bm{s}^{l+1}$, from which the simulated pose $\bm{\hat{x}}^{l+1}$ is derived. By running the policy for $L$ steps, we can obtain the simulated motion sequence $\bm{\hat{x}}^{1:L}$. In implementation, we use reinforcement learning to train the motion imitator $\pi$, where a reward is assigned based on how well the simulated motion $\bm{\hat{x}}^{1:L}$ aligns with input motion $\bm{\tilde{x}}^{1:L}$. In the following, we elaborate on the design of states, rewards, policy and actions.
States. The simulation state $\bm{s}^{l}$ consists of the input motion’s next pose, along with the differences between the input next pose and the current simulated pose across multiple aspects, including joint rotation, position, velocity and angular velocity. This difference information informs the policy of the pose residuals that require compensation, enabling the system to better align the simulated motion with the input motion. Formally, the state is defined as $\bm{s}^{l}=\left[\bm{\tilde{\theta}}^{l+1},\bm{\tilde{p}}^{l+1},\bm{\tilde{\theta}}^{l+1}-\bm{\hat{\theta}}^{l},\bm{\tilde{p}}^{l+1}-\bm{\hat{p}}^{l},\bm{\tilde{v}}^{l+1}-\bm{\hat{v}}^{l},\bm{\tilde{\omega}}^{l}-\bm{\hat{\omega}}^{l}\right]$. Here, $\bm{\theta}$, $\bm{p}$, $\bm{v}$ and $\bm{\omega}$ represent the joint rotation, position, joint velocity, and angular velocity, respectively, $\widetilde{\left[\bm{\cdot}\right]}/\widehat{\left[\bm{\cdot}\right]}$ denote the quantities of the input///simulated motions, and $\left[\bm{\cdot}\right]^{l}$ denotes the quantities at the $l^{th}$ timestep.
Actions. We use the target joint angles of proportional derivative (PD) controllers as the action representation, where the action $\bm{a}^{l}$ specifies the PD target to enable robust motion imitation.
Policy. Following [39, 11, 19], we employ a parameterized Gaussian policy $\pi\left(\bm{a}^{l}|\bm{s}^{l}\right)=\mathcal{N}\left(\mu_{\phi}\left(\bm{s}^{l}\right),\Sigma\right)$, where the mean action $\mu_{\phi}\left(\bm{s}^{l}\right)$ is the output by our motion imitator, a simple multi-layer perceptron network with parameter $\phi$, and $\Sigma$ is a fixed diagonal covariance matrix.
Rewards. The reward function is designed to encourage the simulated motion to match the input motion. At each timestep $l$, the reward $r^{l}$ consists of a mimic reward $r^{l}_{\mathrm{m}}$, a energy penalty $r^{l}_{\mathrm{e}}$ and an adversarial reward $r^{l}_{\mathrm{a}}$, formulating as $r^{l}=r^{l}_{\mathrm{m}}+r^{l}_{\mathrm{e}}+r^{l}_{\mathrm{a}}$. The mimic reward $r^{l}_{\mathrm{m}}$ measures the difference between the simulated pose and input pose across multiple aspects, including joint rotation $\bm{\theta}$, position $\bm{p}$, joint velocity $\bm{v}$ and angular velocity $\bm{\omega}$, as follows:
$r^{l}_{\mathrm{m}}=w_{\theta}\exp\left[{-\alpha_{\theta}\left|\bm{\tilde{\theta}}^{l}-\bm{\hat{\theta}}^{l}\right|}\right]+w_{p}\exp\left[{-\alpha_{p}\left|\bm{\tilde{p}}^{l}-\bm{\hat{p}}^{l}\right|}\right]$ (2)
$+w_{v}\exp\left[{-\alpha_{v}\left|\bm{\tilde{v}}^{l}-\bm{\hat{v}}^{l}\right|}\right]+w_{\omega}\exp\left[{-\alpha_{\omega}\left|\bm{\tilde{\omega}}^{l}-\bm{\hat{\omega}}^{l}\right|}\right],$
[…]
The energy penalty $r^{l}_{\mathrm{e}}$ regulates the policy by discouraging high-frequency foot jitter, a common issue in policy trained without external forces [5]. This energy penalty is computed as:
$r^{l}_{\mathrm{e}}=-0.0005\cdot\left\|\bm{\hat{\nu}}^{l}\bm{\hat{\omega}}^{l}\right\|_{2}^{2},$ (3)
where $\bm{\hat{\nu}}^{l}$ and $\bm{\hat{\omega}}^{l}$ denotes the joint torque and angular velocity of the simulated pose in the $l^{th}$ timestep, and $\left\|\cdot\right\|_{2}$ is the L2 norm.
The adversarial reward $r^{l}_{\mathrm{a}}$ encourages the distribution of simulated motions to align that of input motions through a motion discriminator. Given the state of current simulated pose, denoted as $\bm{\hat{s}}_{\mathrm{a}}^{l}=[\bm{\hat{\theta}}^{l},\bm{\hat{p}}^{l},\bm{\hat{v}}^{l},\bm{\hat{\omega}}^{l}]$, the adversarial reward is calculated as:
$r^{l}_{\mathrm{a}}=-\log\left(1-D_{\psi}\left(\bm{\hat{s}}_{\mathrm{a}}^{l}\right)\right).$ (4)
where $D_{\psi}$ represents the motion discriminator with parameters $\psi$.
Loss. Given the reward $r^{l}$ at each timestep $l$, we train the motion imitator using reinforcement learning (RL). The objective is to maximize the excepted reward, effectively enabling the imitator to closely mimic the input motions. To achieve this, we adopt a standard RL algorithm Proximal Policy Optimization (PPO) [28] to solve for the optimal imitator policy. The objective function is defined as:
$\mathcal{L}_{\text{PPO}}\left(\phi\right)=\mathbb{E}\left[\min\left(\eta(\phi)r^{l},\operatorname{clip}\left(\eta(\phi),1-\gamma,1+\gamma\right)r^{l}\right)\right].$ (5)
Here, $\eta(\phi)$ is the probability ratio $\frac{\pi_{\phi}\left(\bm{a}^{l}\mid\bm{s}^{l}\right)}{\pi_{\phi_{\mathrm{old}}}\left(\bm{a}^{l}\mid\bm{s}^{l}\right)}$, where $\phi$ and $\phi_{\mathrm{old}}$ refer to the parameters of the motion imitator after and before the policy update, $\operatorname{clip}$ is an operation that restricts $\eta(\phi)$ within range $\left[1-\gamma,1+\gamma\right]$ with a hyperparameter $\gamma$.
Motion Discriminator Learning.
We observe that relying solely on physics simulator often produces motions that feel overly mechanical and unnatural. To improve the naturalness of simulated motions, following [22], we introduce a motion discriminator to align the distribution of simulated motions with that of the input motions. Specifically, the motion discriminator $D_{\psi}$ learns to distinguish between the state sampled from input pose ($\bm{\tilde{s}}_{\mathrm{a}}^{l}$) and that of simulated pose ($\bm{\hat{s}}_{\mathrm{a}}^{l}$), with the objective function as:
$\mathcal{L}_{\text{Dis}}\left(\psi\right)$ $=-\mathbb{E}\left[\log\left(D_{\psi}\left(\bm{\tilde{s}}_{\mathrm{a}}^{l}\right)\right)\right]-\mathbb{E}\left[\log\left(1-D_{\psi}\left({\bm{\hat{s}}_{\mathrm{a}}^{l}}\right)\right)\right],$ (6)
where $\psi$ denotes the parameters of the motion distribution.
Initialization and Early Termination.
During training, we employ reference state initialization [21], where a starting point is randomly selected from a motion clip for imitation. We observe that the character frequently falls in the early stages of training. To improve training efficiency and accelerate convergence, we adopt an early stopping strategy. Specifically, we terminate the episode when the MPJPE (Mean Per Joint Position Error) between the state of input pose and that of simulated pose exceeds $0.5$ meter.
Hard Negative Mining.
As training progresses, the motion imitator gradually learns to imitate simple motion sequences. However, more challenging examples in the large-scale motion dataset may be overlooked, limiting the model’s ability to handle difficult samples. To address this, we implement a Hard Negative Mining process that identifies motions where the physical simulator fails to imitate as hard samples. Specifically, a dynamic weight is assigned to each motion sequence in the input data, doubling whenever imitation fails. This process progressively increases the focus on challenging samples, guiding the imitator to effectively learn from difficult examples.
3.3 Motion Generator Fine-tuning
Existing motion generators are often trained with limited real motion data, which may limit their capabilities. However, collecting large-scale real 3D motion data is costly. To overcome this challenge, we leverage a large volume of synthetic, physically-plausible motion data provided by MPR module, to further enhance the capabilities of the motion generator.
As shown in Fig. 2, in the second stage, we further finetune the motion generator using the physics-refined motions produced by the MPR module. Notably, since the physical simulator cannot replicate non-grounded motions (e.g., sitting on a chair or swimming), such simulated motions may deviate from the true data distribution. To this end, we apply an Imitation Selection Operation to filter out simulated data of non-grounded motions. Specifically, we calculate the average per-joint position error (MPJPE) between the samples before and after physical optimization. A threshold $\tau$ is set to determine whether to accept the physically refined motion $\bm{\hat{x}}^{1:L}$ (with $\text{MPJPE}<\tau$) or input motions $\bm{\tilde{x}}^{1:L}$ (with $\text{MPJPE}>\tau$). The selected data is then paired with the original condition signals (e.g., text or music). Ultimately, a large-scale, physically plausible motion data is constructed for fine-tuning the motion generator.
Following [7, 32], we use the mean squared error to optimize the motion generator. Denote the selected motion sequence paired with the condition as $\left(\bm{x}^{1:L},c\right)$, the objective function for motion generator $f_{\xi}$ is defined as:
$\mathcal{L}_{\text{MG}}\left(\xi\right)=\mathbb{E}\left[\left\|\bm{x}^{1:L}-f_{\xi}\left(c\right)\right\|_{2}^{2}\right]$ (7)
where $\xi$ is the parameters of the motion generator.
Inference.
As shown in Fig 2, after the two-stage optimization, the finetuned motion generator (Stage 2) and the trained MPR module (Stage 1) are combined to perform inference. To mitigate simulation error in non-grounded motions, the imitation selection operation described above is used as a post-process step.
4 Experiment
Extensive experiments evaluate the performance of our Morph across multiple motion generation tasks and datasets. Specifically, we assess Morph on two motion generation tasks, text-to-motion and music-to-dance. As Morph is agnostic to specific instantiation of generation models, we combine Morph with three types of generation models for text-to-motion, i.e., diffusion-based models, including MotionDiffuse [40] and MDM [31], autoregressive-based models, including T2M-GPT [40] and generative make modeling, including MoMask [7]. For music-to-dance, we combine Morph with the diffusion-based model EDGE [33] and autoregressive model Bailando [29]. The Appendix provides additional visualizations and experimental results showcasing Morph’s performance on varying amounts of noisy motion data, different threshold $\tau$ in imitation selection opeartion, and multi-round optimization of the MPR module and motion generator.
Table 1: Ablation study on Morph-MoMask (combined with MoMask [7] generator) for text-to-motion task on HumanML3D dataset. IS: imitation selection operation; Adversarial: using adversarial reward training MPR module; Energy: using energy reward training MPR module; Real Data: using real motion data training MPR module; FT: fine-tuning Motion Generator with physics-refined motions. The arrows ($\uparrow$/$\downarrow$ / ↓) indicate that higher/smaller values are better.
Model Number Methods Common Generation Metrics Physical Plausibility Metrics
IS Adversarial Energy Real Data FT RTOP-3 $\uparrow$ FID $\downarrow$ PFC $\downarrow$ Penetrate $\downarrow$ Float $\downarrow$ Skate $\downarrow$ IFR $\downarrow$
$\mathrm{A}$ Baseline (only motion generator MoMask [7]) 0.807 0.045 1.058 23.152 10.660 5.262 -
$\mathrm{B}$ 0.792 0.194 0.852 0.000 2.272 0.020 0.0155
$\mathrm{C}$ 0.782 0.276 0.735 0.000 2.554 0.032 0.0362
$\mathrm{D}$ 0.790 0.165 0.715 0.000 2.376 0.026 0.0272
$\mathrm{E}$ 0.801 0.077 0.672 0.000 2.271 0.011 0.0155
$\mathrm{F}$ 0.785 0.183 0.749 0.000 2.451 0.017 0.0338
$\mathrm{G}$ 0.816 0.043 0.651 0.000 2.146 0.010 0.0151
4.1 Datasets and Evaluation Metrics
HumanML3D. HumanML3D [9] is a large-scale 3D human motion-language dataset, consisting of 14,616 motion clips and 44,970 motion descriptions annotations. The total motion duration is $28.59$ hours, with each motion clip downsampled to 20 FPS and accompanied by 3-4 textual descriptions. The dataset is split into training, validation and test sets in an $80\%$, $5\%$ and $15\%$ ratios.
AIST++. AIST++ [15] is a human dance-music dataset, which includes $992$ pieces of high-quality dance sequence across ten dance genres, each dance paired with a corresponding music. The dance durations range from $7.4$ to $48$ seconds, with a frame rate of 60 FPS. The dataset is split into training and evaluation sets, with allocations of $952$ and $40$, respectively.
Evaluation Metrics. We evaluate motion generation from two perspectives: common generation metrics and physical plausibility metrics. We first present the common generation metrics. For text-to-motion generation, following [8, 41], we use three standard metrics: Frechet Inception Distance (FID) that measures the distance between the generated and ground-truth motion distributions; R-Precition (including RTOP-1, RTOP2, and RTOP-3) that computes the retrieval accuracy of the generated motions with respect to the input text; Diversity (Div) that measures the diversity of generated motions. For music-to-dance generation, following [16, 38, 29], we use five standard metrics: FIDk and Divk that calculate Frechet Inception Distance and Diversity on kinetic features (denoted as ‘$k$’); FIDg and Divg that calculate Frechet Inception Distance and Diversity on geometric features (denoted as ‘$g$’); Beat Align Score (BAS) that measures the alignment between input music and generated dances.
Next, we present the physical plausibility metrics. Following [39, 33], we use four physic-based metrics to assess the physical plausibility of generated motions: Penetrate that measures ground penetration; Float that measures floating; Skate that measures foot sliding; Physical Foot Contact score (PFC) that measures the realism of foot-ground contact. Additionally, we introduce an imitation selection metric, imitation failure rate (IFR), which calculates the failure rate of the physics refinement module in imitating motion.
4.2 Implementation Details
Training Setup. We implement Morph based on PyTorch. The NVIDIA’s Isaac Gym is used as the physics simulator. Rather than using real motion data, we train Morph on synthetic motion data produced by a pre-trained model, increasing the motion data volume to three times the original. In the training phase of MPR module (Stage 1), we use the Adam optimizer with a batch size of $64$ and a learning rate of $4\times 10^{-5}$. During the finetuning of motion generator (Stage 2), we follow the original training setup, only adjusting the learning rate to $1\times 10^{-5}$. For hyper-parameters setting, we set $w_{\theta}$, $w_{q}$, $w_{v}$ and $w_{w}$ (Eq. 2) to $0.5$, $0.3$, $0.1$ and $0.1$, $\alpha_{\theta}$, $\alpha_{q}$, $\alpha_{v}$ and $\alpha_{w}$ (Eq. 2) to $100$, $10$, $0.1$ and $0.1$, $\gamma$ (Eq. 5) is set to $0.1$ and the imitation selection threshold $\tau$ is set to $0.5$. All our experiments are conducted on 8 Tesla V100 GPUs. Further details on implementation and hyperparameter analysis are provided in the Appendix.
Data Preprocessing. In text-to-motion generation, some motion sequences involve environmental interactions, such as sitting down, climbing stairs, and swimming, which cannot be simulated by current physical simulator. Therefore, we annotate each text description in HumanML3D [9] based on its semantic content. Text Descriptions involving interaction-based motions are labeled as ‘00’, while other descriptions are labeled as ‘$1$’. Text labeled as ‘$1$’ will be used to train the MPR module in Stage 1. We will release this annotation on HumanML3D in the final version.
Table 2: Comparison results for text-to-motion task on HumanML3D dataset. Morph is combined with different types of motion generators. MG: Motion Generator; MPR: Motion Physics Refinement module; FT: fine-tuning motion generator with the physics-refined motion data. $\dagger$ denotes Morph without fine-tuning the motion generator (only Stage 1 training)
MG MPR FT Common Generation Metrics Physical Plausibility Metrics
RTOP-1 $\uparrow$ RTOP-3 $\uparrow$ FID $\downarrow$ Diversity $\uparrow$ PFC $\downarrow$ Penetrate $\downarrow$ Float $\downarrow$ Skate $\downarrow$ IFR $\downarrow$
PhysDiff w/ MD [39] - - - 0.780 0.551 - - 0.898 1.368 0.423 -
PhysDiff w/ MDM [39] - - - 0.631 0.433 - - 0.998 2.601 0.512 -
Reindiffuse [11] - - - 0.622 0.385 - - 0.000 0.711 0.058 -
MDM [31] - - 0.455 0.749 0.489 9.920 0.811 17.384 17.502 3.540 -
[…]
Morph-MDM 0.467 0.752 0.485 9.863 0.707 0.000 2.261 0.018 0.0154
MotionDiffuse(MD) [40] - - 0.491 0.782 0.630 9.410 0.533 16.670 5.698 3.419 -
[…]
Morph-MD 0.494 0.786 0.571 9.492 0.416 0.000 0.856 0.031 0.0166
T2M-GPT [40] - - 0.491 0.775 0.116 9.761 0.998 72.250 8.918 7.801 -
[…]
Morph-T2M-GPT 0.497 0.781 0.109 9.768 0.751 0.000 2.703 0.042 0.0197
MoMask [7] - - 0.521 0.807 0.045 9.641 1.058 23.152 10.660 5.262 -
[…]
Morph-MoMask 0.525 0.816 0.043 9.687 0.651 0.000 2.146 0.010 0.0151
Table 3: Comparison results on common generation metrics for text-to-motion on HumanML3D dataset.
Methods RTOP-1 $\uparrow$ RTOP-2 $\uparrow$ RTOP-3 $\uparrow$ FID $\downarrow$ MM-Dist $\downarrow$ Diversity $\uparrow$ MModality $\uparrow$
MDM [31] 0.455 0.645 0.749 0.489 3.330 9.920 2.290
MotionDiffuse (MD) [41] 0.491 0.681 0.782 0.630 3.113 9.410 1.553
MLD [3] 0.481 0.673 0.772 0.473 3.196 9.724 2.413
T2M-GPT [40] 0.491 0.680 0.775 0.116 3.118 9.761 1.856
AttT2M [47] 0.499 0.690 0.786 0.112 3.038 9.700 2.452
MMM [24] 0.515 0.708 0.804 0.089 2.926 9.577 1.226
MoMask [7] 0.521 0.713 0.807 0.045 2.958 9.641 1.241
BAMM [23] 0.525 0.720 0.814 0.055 2.919 9.717 1.687
Morph-MDM 0.467 0.658 0.752 0.485 3.106 9.863 2.177
Morph-MD 0.494 0.687 0.786 0.571 3.001 9.492 1.584
Morph-T2M-GPT 0.497 0.684 0.781 0.109 2.997 9.768 1.903
Morph-MoMask 0.525 0.722 0.816 0.043 2.923 9.687 1.652
Table 4: Comparison results for music-to-dance on AIST++ dataset. $\dagger$ denotes Morph without fine-tuning motion generator.
MG MPR FT Common Generation Metrics Physical Plausibility Metrics
FIDk $\downarrow$ FIDg $\downarrow$ Divk $\uparrow$ Divg $\uparrow$ BAS $\uparrow$ PFC $\downarrow$ Penetrate $\downarrow$ Float $\downarrow$ Skate $\downarrow$ IFR $\downarrow$
FACT [14] - - 35.35 22.11 5.94 6.18 0.2209 - - - - -
TM2D [6] - - 19.01 20.09 9.45 6.36 0.2049 - - - - -
EDGE [33] - - 42.16 22.12 3.96 4.61 0.2334 0.610 76.490 60.982 7.906 -
Bailando [29] - - 28.16 9.62 7.83 6.34 0.2332 0.074 31.183 12.650 4.466 -
[…]
Morph-EDGE 39.29 19.88 4.91 5.23 0.2327 0.287 0.000 3.524 0.010 0.0085
[…]
Morph-Bailando 26.32 9.43 7.86 6.55 0.2411 0.045 0.000 2.061 0.023 0.0059
The generated motion sequences may exhibit issues such as body leaning, floating and ground penetration. When imported into the engine, these issues can cause instability in the robot, leading to falls, bouncing off the ground, or dropping from mid-air. To address this, we apply a preprocessing step to the motion sequences. Specifically, we first compute the body’s tilt angle, defined as the angle between the projection of the center of mass onto the ground and the line connecting both feet, and apply this angle to the pelvis for the entire sequence. To correct for floating and penetration, we determine the lowest point height and adjust the entire sequence by this offset. The resulting sequence is then used for training and inference. More details can be found in the Appendix.
4.3 Ablation Studies
In this section, we conduct ablation studies to validate the effectiveness of each part of our method. We use MoMask [7] as the motion generator in Morph, denoted as Morph-MoMask, throughout the ablation studies. The ablation results are shown in Tab. 1.
Effectiveness of training MPR module using only generated data. We investigate the effectiveness of using only synthetic motion data from two perspectives: its overall effectiveness and its performance relative to training with real motion data. First, we compare the baseline model $\mathrm{A}$ (only the MoMask generator) and model $\mathrm{E}$ which combines MoMask with the trained MPR module. As shown in Tab. 1, model $\mathrm{E}$ outperforms model $\mathrm{A}$ on all physical metrics ($0.651/1.058$ on PFC, $0.0/23.152$ on Penetrate, $2.146/10.66$ on Float, $0.01/5.262$ on Skate), while maintaining competitive performance on generation metrics. These results confirm the feasibility of training the MPR module solely with synthetic data. Next, we compare model $\mathrm{E}$ with model $\mathrm{F}$, which uses real motion data to train the MPR module. As shown in Tab. 1, model $\mathrm{F}$ achieves inferior performance compared to model $\mathrm{E}$. This can be attributed to the significant domain gap between the training data (real motions) and test data (generated motions). This gap hinders the MPR module trained on real data from effectively adapting to the generated motion data, leading to performance decline.
Effectiveness of adversarial reward and energy reward. When training the MPR module, we add an adversarial reward and energy reward in addition to the common imitation task reward. In Tab. 1, comparing model $\mathrm{E}$ with model $\mathrm{C}$, which omits the adversarial reward, we observe a performance decline in both generation and physical metrics on model $\mathrm{C}$. The significant drop in FID metric indicates that, without the adversarial reward, the simulated motions lack the distribution constraints of the input motions, leading to a deviation from the intended motion distribution. Next, we compare model $\mathrm{E}$ with $\mathrm{D}$, which omits the energy reward. As shown in Tab. 1, model $\mathrm{E}$ outperforms model $\mathrm{D}$ in generation metrics. We argue that introducing energy reward helps suppress high-frequency jitter during the humanoid robot control process, resulting in more natural motions.
Effectiveness of imitation selection strategy. Like [39, 11], the proposed Morph focuses on addressing physical inconsistencies in non-interactive and ground-contact human motions. In training and testing data, there are some interactive and non-grounded motions, such as sitting on a chair, swimming, or climbing stairs. To ensure the quality of the motions after refining, we design an imitation selection operation to filter out the simulations of such motions. In Tab. 1, we compare model $\mathrm{E}$ with $\mathrm{B}$, which omits the imitation selection operation. As shown, model $\mathrm{E}$ outperforms $\mathrm{B}$ in both generation and physic metrics, validating the effectiveness of the imitation selection strategy.
Effectiveness of finetuning motion generator using simulated motions. Furthermore, we compare model $\mathrm{E}$ and model $\mathrm{G}$, which finetunes the motion generator using the physically refined motions produced by the MPR module. As shown in Tab. 1, model $\mathrm{G}$ achieves further gains in both generation and physic metrics, with RTOP-3 increasing by $1.5\%$ and FID improving by $0.034$. These results suggest that the proposed MPR module can, in turn, enhance the performance of the motion generator.
4.4 Evaluation with Different Motion Generators
We evaluate the adaptability of our Morph with different motion generators, including MDM [31], Motiondiffuse [41], T2M-GPT [40], and MoMask [7]. As shown in Tab. 2, Morph significantly enhances physical fidelity while maintaining competitive generation metrics across different motion generation models. For example, after integrating Morph with MDM (Morph-MDM), RTOP-1 metric improved by $0.012$. In terms of physical plausibility metrics, penetration dropped to zero, float decreased from $17.502$ to $2.261$, and skate reduced from $3.540$ to $0.018$. Similar performance improvements are observed with other motion generators. The consistent gains across different generators highlight the versatility of our Morph framework.
Figure 3: Qualitative comparison between our Morph-MoMask and MoMask in text-to-motion task. Morph-MoMask significantly reduces physical artifacts such as leaning forward, floating and penetration.
4.5 Comparisons with State-of-the-arts
Main results on Text-to-Motion Generation. On the text-to-motion dataset, HumanML3D, we compare Morph framework with other state-of-the-art methods. The comparison results are shown in Tab. 2 and Tab. 3. As shown, Morph achieves significant gains in physical metrics while maintaining competitive performance in generation metrics, demonstrating its capability to enhance physical plausibility. Fig. 3 illustrates the generated motions from both our Morph-MoMask and MoMask. As shown, motions generated by MoMask often exhibit physically unrealistic artifacts, such as penetration, floating, and leaning forward. In contrast, Morph-MoMask effectively reduces these artifacts, producing motions that are both physically plausible and realistic.
Main results on Music-to-Dance Generation. On the music-to-dance dataset, AIST++, we compare Morph combined with EDGE [33] and Bailando [29] generators against other state-of-the-art methods. As shown in Tab. 4, Morph significantly improves physical metrics compared to existing methods. In terms of generation metrics, our Morph-EDGE model achieves the best results in multiple metrics, including FIDg, Divg, and BAS, while also demonstrating competitive performance in FIDk metric. These results validate the superiority of our framework in music-to-dance generation.
5 Conclusion
In this paper, we present Morph, a model-agnostic physical optimization framework, designed to enhance physical plausibility in motion generation without relying on costly real-world motion data. To accomplish this, we first leverage a pretrained motion generator to synthesize large-scale noisy motion data. We then introduce a Motion Physics Refinement module, which utilizes these synthesized data to train a motion imitator that enforces physical constraints. The physically refined motions can, in turn, be used to finetune the motion generator, further enhancing its capabilities. Our framework is compatible with various motion generation models across both text-to-motion and music-to-dance generation tasks. Extensive experimental results demonstrate that Morph substantially improves physical plausibility while achieving competitive generation quality.
Appendix
In this Appendix, we present more details for Morph, including data preprocess, additional experimental results, qualitative comparisons. First, we describe the data preprocessing procedure used for training the Motion Physics Refinement (MPR) module with generated motion data (Sec. A). Then, we present experimental results analyzing the impact of $\tau$ in the imitation selection operation (Sec. B), the effect of varying the quantity of noisy motion data for MPR training (Sec. C), and effect of the number of training rounds for Morph (Sec. D). Finally, we provide additional qualitative comparisons for text-to-motion and music-to-dance tasks (Sec. E).
Appendix A Details for Data Preprocess
As discussed in the main text, the generated motion sequences may exhibit issues such as body leaning, floating and ground penetration. When imported into the simulator, these issues can cause instability in the robot, potentially causing it to fall, bounce off the ground, or drop from mid-air. To address this issue, we apply a preprocessing step to the motion sequences, detailed in Fig. 4 and Alg. 1. Specifically, we first compute the body’s tilt angle, defined as the angle between the projection of the center of mass onto the ground and the line connecting both feet. If this angle exceeds $10^{\circ}$, we apply the necessary adjustment to the pelvis throughout the sequence. To correct floating and penetration, we determine the lowest mesh height and adjust the entire sequence by this offset. The preprocessed sequence is then used for training and inference.
Figure 4: A flowchart illustrating the data preprocessing process. The parameters are calculated from the first frame and then applied to all generated motion sequences before they are fed into the MPR module.
Algorithm 1 Preprocessing Motion Sequences
1:Motion sequence $S$ with frames $F_{1},F_{2},\ldots,F_{n}$
2:Preprocessed motion sequence $S^{\prime}$
3:Step 1: Calculate the angle $\theta$
  1. (1)
    Compute the projection of the center of mass of $F_{1}$ onto the ground.
  2. (2)
    Determine the line connecting the pelvis point and the center of both feet in $F_{1}$.
  3. (3)
    Calculate the angle $\theta$ between the projection and the line.
4:Step 2: Correct posture if $\theta>10^{\circ}$
  1. (1)
    Apply an additional rotation to the pelvis for the entire sequence $S$.
5:Step 3: Ensure $F_{1}$ is on the ground
  1. (1)
    Infer the lowest point height $h$ of the mesh in $F_{1}$.
  2. (2)
    Add a uniform offset to the entire sequence $S$.
6:Step 4: Output the preprocessed sequence $S^{\prime}$.
[…]
Methods Common Generation Metrics Physical Plausibility Metrics
RTOP-1 $\uparrow$ RTOP-3 $\uparrow$ FID $\downarrow$ Diversity $\uparrow$ PFC $\downarrow$ Penetrate $\downarrow$ Float $\downarrow$ Skate $\downarrow$ IFR $\downarrow$
$\tau$=0.0 0.521 0.807 0.045 9.641 1.058 23.152 10.660 5.262 -
$\tau$=0.1 0.520 0.806 0.051 9.633 0.879 3.567 4.783 2.132 0.1287
$\tau$=0.2 0.518 0.806 0.058 9.629 0.774 0.842 4.020 1.061 0.0543
$\tau$=0.3 0.515 0.804 0.071 9.578 0.760 0.058 3.202 0.533 0.0262
$\tau$=0.4 0.514 0.803 0.073 9.582 0.728 0.003 2.998 0.217 0.0162
$\tau$=0.5 0.511 0.801 0.077 9.579 0.672 0.000 2.271 0.011 0.0155
$\tau$=0.6 0.508 0.799 0.080 9.575 0.686 0.000 2.268 0.011 0.0147
$\tau$=0.7 0.506 0.799 0.081 9.543 0.663 0.000 2.102 0.008 0.0131
$\tau$=0.8 0.500 0.795 0.081 9.522 0.647 0.000 2.028 0.006 0.0126
$\tau$=0.9 0.496 0.793 0.087 9.412 0.638 0.000 1.989 0.005 0.0120
$\tau$=1.0 0.488 0.790 0.088 9.260 0.625 0.000 1.984 0.005 0.0113
[…]
Methods Common Generation Metrics Physical Plausibility Metrics
RTOP-1 $\uparrow$ RTOP-3 $\uparrow$ FID $\downarrow$ Diversity $\uparrow$ PFC $\downarrow$ Penetrate $\downarrow$ Float $\downarrow$ Skate $\downarrow$ IFR $\downarrow$
[…]
Table 7: Comparison of text-to-motion with multi-round optimization of the MPR module and motion generator based on Morph-MoMask. We set $\tau$ as 0.5 and use the total number of generated noisy motion data to train.
Methods Common Generation Metrics Physical Plausibility Metrics
RTOP-1 $\uparrow$ RTOP-3 $\uparrow$ FID $\downarrow$ Diversity $\uparrow$ PFC $\downarrow$ Penetrate $\downarrow$ Float $\downarrow$ Skate $\downarrow$ IFR $\downarrow$
One-Round w/o FT 0.511 0.801 0.077 9.579 0.672 0.000 2.271 0.011 0.0155
One-Round 0.525 0.816 0.043 9.687 0.651 0.000 2.146 0.010 0.0151
Two-Round w/o FT 0.525 0.818 0.043 9.689 0.635 0.000 2.132 0.010 0.0137
Two-Round 0.527 0.821 0.044 9.692 0.631 0.000 2.117 0.008 0.0134
Appendix B Effect of $\tau$ in Imitation Selection on Morph
[…]
Appendix C Effect of Varying Amounts of Noisy Motion Data on Morph
[…]
Appendix D Effect of Multi-Round Optimization of the MPR module and MG on Morph
In Tab. 7, we analyze the effect of multi-round optimization of the Physics Refinement (MPR) module and Motion Generator (MG) on Morph using Morph-MoMask. To further validate the effectiveness of this round-based training approach in enhancing both the MG and the MPR module, we conducted an additional round of training beyond this single-round training described in the main text. This extra round explores the potential for mutual enhancement between the two modules. In Tab. 7, the following terms are defined:
  • One-Round w/o FT: The first round of training where only the MPR module is trained.
  • One-Round: The first round of training that includes both training the MPR module and fine-tuning the MG.
  • Two-Round w/o FT: Training the MPR module again using the motion data generated by the fine-tuned MG from the first round.
  • Two-Round: Fine-tuning the Motion Generator using the results from Two-Round w/o FT.
As shown in Tab. 7, in the first round of training, MG improves the performance of MPR module, enhancing the physical quality of its generated motion. The refined motion data from the trained MPR module is then used to fine-tune the MG, boosting its performance further. In the second round, the fine-tuned MG from the first round is used to generate training data for the MPR module (initialized with first-round weights). We observed improvements in Two-Round w/o FT compared to One-Round, with PFC increasing by 0.016, Float by 0.014, and IFR decreasing, indicating enhanced motion imitation by the MPR module. After fine-tuning the MG once again, Two-Round shows improvements in the RTOP-1 and RTOP-3 metrics. These results clearly demonstrate that the MG and MPR modules can mutually enhance each other. Moreover, alternating training between the MG and MPR modules across multiple rounds can further improve the performance of Morph.
Appendix E More Qualitative Results
Fig. 5 and Fig. 6 provide the additional qualitative results for the text-to-motion and music-to-dance generation tasks using Morph.
As shown in Fig. 5, in the text-to-motion generation task, floating and penetration are common artifacts in motion generation, often resulting from inaccuracies in the estimation of translation. However, Morph effectively addresses these issues, successfully mimicking the input motion and demonstrating a significant improvement in mitigating these artifacts. The generated motions are both physically plausible and realistic, showcasing Morph’s enhanced performance in this task.
As shown in Fig. 6, in the music-to-dance generation task, floating and penetration are the most prominent issues. Due to the faster frequency of dance movements, these artifacts occur more frequently. Morph effectively mitigates these issues, generating motions that are not only physically plausible but also exhibit a higher degree of realism.
In summary, Morph demonstrates significant improvements in both the text-to-motion and music-to-dance tasks. By accurately estimating translational motion, Morph is able to generate motions that are not only physically feasible but also exhibit a higher degree of realism.
Figure 5: Qualitative comparisons for text-to-motion on HumanML3D test set between Morph-MoMask and MoMask.
Figure 6: Qualitative comparisons for music-to-dance on AIST++ test set between Morph-Bailando and Bailando. For music-to-dance, the testing music samples will be used as inputs.

### v3

@@
 Abstract
-Human motion generation plays a vital role in applications such as digital humans and humanoid robot control.
-However, most existing approaches disregard physics constraints, leading to the frequent production of physically implausible motions with pronounced artifacts such as floating and foot sliding.
-In this paper, we propose Morph, a Motion-free physics optimization framework, comprising a Motion Generator and a Motion Physics Refinement module, for enhancing physical plausibility without relying on costly real-world motion data.
-Specifically, the Motion Generator is responsible for providing large-scale synthetic motion data, while the Motion Physics Refinement Module utilizes these synthetic data to train a motion imitator within a physics simulator, enforcing physical constraints to project the noisy motions into a physically-plausible space.
-These physically refined motions, in turn, are used to fine-tune the Motion Generator, further enhancing its capability.
-Experiments on both text-to-motion and music-to-dance generation tasks demonstrate that our framework achieves state-of-the-art motion generation quality while improving physical plausibility drastically.
+Human motion generation has been widely studied due to its crucial role in areas such as digital humans and humanoid robot control.
+However, many current motion generation approaches disregard physics constraints, frequently resulting in physically implausible motions with pronounced artifacts such as floating and foot sliding.
+Meanwhile, training an effective motion physics optimizer with noisy motion data remains largely unexplored.
+In this paper, we propose Morph, a Motion-Free physics optimization framework, consisting of a Motion Generator and a Motion Physics Refinement module, for enhancing physical plausibility without relying on expensive real-world motion data.
+Specifically, the motion generator is responsible for providing large-scale synthetic, noisy motion data, while the motion physics refinement module utilizes these synthetic data to learn a motion imitator within a physics simulator, enforcing physical constraints to project the noisy motions into a physically-plausible space.
+Additionally, we introduce a prior reward module to enhance the stability of the physics optimization process and generate smoother and more stable motions.
+These physically refined motions are then used to fine-tune the motion generator, further enhancing its capability.
+This collaborative training paradigm enables mutual enhancement between the motion generator and the motion physics refinement module, significantly improving practicality and robustness in real-world applications.
+Experiments on both text-to-motion and music-to-dance generation tasks demonstrate that our framework achieves state-of-the-art motion quality while improving physical plausibility drastically.
+Figure 1: Examples of physical inconsistencies in generations.
 1 Introduction
-Accurate human motion generation is pivotal in various applications, including robotics, video gaming and virtual reality [10, 36, 45, 44, 1].
-Recent advances in AI have paved the way for novel approaches to motion generation, enabling various control conditions including textual descriptions and music pieces.
-Mainstream approaches can be divided into three categories: conditional diffusion models [32, 42, 39, 33, 16], conditional autoregressive models [23, 13, 46, 47, 18] and generative masked modeling [7, 24].
-These motion generation models have made significant progress in capturing the complex multimodal distribution of human motions, greatly improving motion generation quality.
+Accurate human motion synthesis plays a crucial role in applications such as robotics, video gaming, and virtual reality [9, 38, 47, 46, 1].
+Recent advances in AI have enabled motion generation under various control conditions, including textual descriptions and music inputs.
+Current approaches can be categorized into three main paradigms: conditional diffusion models [34, 44, 41, 35, 15], conditional autoregressive models [26, 12, 48, 49, 19], and generative masked modeling [6, 27].
+These models have significantly improved the quality of generated motion by better capturing the complex multimodal distribution of human movement.
 However, most existing motion generation approaches overlook a fundamental aspect of human motion: the laws of physical.
@@
 Given human sensitivity to even slight physical inconsistencies, these physically implausible motions hinder many real-world applications such as animation and virtual reality [41, 30, 11].
-Figure 1: Examples of physical inconsistencies in generations.
-Recently, a few studies [39, 11] have attempted to enhance the physical plausibility of motion generation.
-For example, PhysDiff [39] and Reindiffuse [11] incorporate physical constraints into the denosing diffusion process.
-However, these approaches face several limitations.
-First, these works [39, 11] customize physics optimization operation for a specific type of generation model, i.e., text-conditional diffusion model.
-This customization restricts their transferability to a broader range of generation models (such as autoregressive-based models) and motion generation tasks (such as music-to-dance).
-Second, these works apply physics and diffusion iteratively, embedding physics optimization into multiple steps of diffusion process.
+Recently, a few studies [41, 10] have attempted to enhance the physical plausibility of motion generation, making the generated motions more realistic.
+For example, PhysDiff [41] and Reindiffuse [10] incorporate physical constraints into the denoising diffusion process.
+However, these approaches face two main limitations.
+On the one hand, although these works [41, 10] introduce promising approaches for motion physics optimization, their generalizability and transferability across different types of motion generation models and tasks have not been fully demonstrated or explored.
+On the other hand, these works apply physics and diffusion iteratively, embedding physics optimization into multiple steps of diffusion process.
 The frequent execution of physics optimization significantly increases computational costs during inference.
-Last but most importantly, these methods require large-scale, high-quality realistic motion data to train the physics optimization mechanism.
-However, collecting such realistic motion data is challenging.
+Additionally, prior studies [22, 21, 20] have explored physics-based human motion imitation using a physics simulator, typically relying on large-scale real 3D motion data as priors to enhance accuracy and naturalness in imitation learning.
+However, collecting large-scale realistic motion data is challenging.
 A natural question arises: Is it possible to learn an efficient, model-agnostic physical optimizer without relying on real motion data?
@@
 Morph employs a two-stage optimization process.
-In the first stage, using large-scale synthetic noisy motion data produced by the motion generator, MPR module is optimized to project input motions into a physically-plausible space.
-Specifically, the MPR module designs a motion imitator that controls a character agent to mimic the given noisy motions within a physics simulator.
+In the first stage, using large-scale noisy motion data produced by the motion generator, MPR module is optimized to project input motions into a physically-plausible space.
+Specifically, MPR module designs a motion imitator that controls a character agent to mimic the given noisy motions within a physics simulator.
 The simulator enforces multiple physical constraints, effectively reducing artifacts such as floating and foot sliding.
-To further ensure the naturalness of simulated motions, MPR module introduces a motion discriminator to align the distribution of physics-refined motions with that of input motions.
-Feedback signals from both physics simulator and motion discriminator guide the optimization of the motion imitator via reinforcement learning.
-In the second stage, we observe that due to the high cost of collecting 3D motion data, existing generators typically rely on limited motion data for training, which restricts their generative capabilities.
-In this stage, by leveraging the trained MPR module, a large-scale set of physically plausible, high-quality motion data can be created to finetune the generator, further enhancing its ability to generate realistic motion.
-During inference, the finetuned motion generator and MPR module work in tandem to generate physics-plausible and high-quality motions.
-Through above process, Morph effectively decouples physic optimization from generation model, requiring only a single step of physical refinement and without relying on real motion data.
-This results in an economical, efficient, and model-agnostic physical optimizer.
-We evaluate our Morph framework on two motion generation tasks: text-to-motion and music-to-dance generation.
-Since Morph is agnostic to specific instantiation of generation models, we test it with three types of generation models, i.e., diffusion-based, autoregressive-based and generative mask modeling.
-Extensive experiments demonstrate that Morph achieves significant improvements in physical error metrics, while also achieving competitive generation metrics across different generators and tasks, despite not being trained on real motion data.
-Figure 2: An overview of the Morph framework.
-Morph comprises a Motion Generator and a Motion Physics Refinement module.
-Morph employs a two-stage training process: Motion Physics Refinement module training and Motion Generator fine-tuning.
-And a Imitation Selection Operation is employed to ensure the motion quality after physics refinement.
+To further ensure the naturalness of simulated motions, MPR module introduces a prior reward module to align the distribution of physics-refined motions with that of input motions.
+This prior reward module can provide smoother and more stable rewards.
+Feedback signals from both physics simulator and prior reward module guide the optimization of the motion imitator via reinforcement learning.
+In the second stage, we address the limitations of pretrained motion generators, which may not be trained on data that adheres to physical constraints and thus produce suboptimal results.
+To this end, we finetune the motion generator using the physic-refined motion data produced by MPR module, further enhancing its generation capability.
+During inference, the fine-tuned motion generator and MPR module work in tandem to generate physics-plausible and high-quality motions.
+By fostering the synergy between motion generation and motion physics refinement, morph provides a scalable, efficient, and generalizable solution to realistic human motion synthesis.
+We evaluate our morph framework on text-to-motion and music-to-dance tasks, using diffusion-based, autoregressive, and generative mask models.
+Extensive experiments demonstrate that morph achieves significant reduces physical errors, while maintaining competitive motion quality across different generators and tasks, despite not being trained on real motion data.
+Our contributions can be summarized as follows:
+  • We introduce Morph, a generalizable framework that enhances physical plausibility and generation quality across various motion generation models and tasks.
+Also, a prior reward module is designed to accelerate imitation learning and improve motion realism in physics simulation.
+  • We design a collaborative training paradigm, where the Motion Generator provides large-scale synthetic motion data to enhance the Motion Physics Refinement module, which, in turn, refines motion quality by feeding back physics-refined data.
+  • Extensive qualitative and quantitative experiments demonstrate that Morph achieves competitive physical plausibility and generation performance.
 2 Related Work
 Human Motion Generation.
-Motion generation is a long-history task that can be conditioned on various signals, such as text description, music and action [48, 7, 40, 8, 29, 33, 37, 2, 4, 30, 43].
+Motion generation is a long-history task that can be conditioned on various signals, such as text description, music and action [50, 6, 42, 7, 31, 35, 39, 16, 17, 2, 4, 32, 45].
 Our work specifically focuses on text-to-motion [7, 42, 43, 34, 6, 27] and music-to-dance generation [35, 31, 15, 23].
@@
 In contrast, our proposed MPR module effectively enforces physical constraints, significantly enhancing the physically plausibility of generated motions.
+Figure 2: An overview of the Morph framework.
+Morph comprises a Motion Generator and a Motion Physics Refinement module.
+Morph employs a two-stage training process: Motion Physics Refinement module training and Motion Generator fine-tuning.
+And a Imitation Selection Operation is employed to ensure the motion quality after physics refinement.
+The solid curved arrows on the left and right (in orange and green) represent the iterative, collaborative optimization between Stage 1 and Stage 2.
 Physically Plausible Motion Generation.
@@
 EDGE [35] introduces auxiliary losses to align specific aspects of physical realism, specifically for generating more physically plausible dances.
-However, these methods tailor physic optimization to diffusion models and require real motion data, limiting their flexibility and scalability.
-In contrast, we propose a model-agnostic, motion-free physics optimization framework that decouples physics optimization from the generation model and learns to enhance physical plausibility without real motion data, offering a more economical and versatile solution.
+However, these methods rely on real motion data and have not fully demonstrated generalizability across different motion generation models and tasks, limiting their flexibility and scalability.
 3 Method
@@
 In the second stage, these physics-refined motions, in turn, are used to fine-tune Motion Generator, further enhancing its capabilities (Sec 3.3).
-Through this two-stage training process, Morph drastically reduces physical errors while improving the overall quality of generated motions.
+Through these two-stage training process, morph drastically reduces physical errors while improving the overall quality of generated motions.
 3.1 Noisy Motion Data Generation
@@
 As shown in Fig.
-2, the MPR module consists of three components: a motion imitator, a physics simulator and a motion discriminator.
+2, the MPR module consists of three components: a motion imitator, a physics simulator and a motion prior reward model.
 Specifically, the motion imitator controls a simulated character to mimic the input motion $\bm{\tilde{x}}^{1:L}$ within the physics simulator.
 The resulting motion $\bm{\hat{x}}^{1:L}$ from the physics simulator is considered physically plausible, as it adheres to the laws of physics.
-Additionally, the motion discriminator seeks to align the distribution of refined motions $\bm{\hat{x}}^{1:L}$ with that of input motions $\bm{\tilde{x}}^{1:L}$, further ensuring the naturalness and realism of the physics-refined motions.
+Additionally, the prior reward model penalize the non-human motions further ensuring the naturalness and realism of the physics-refined motions.
 Motion Imitator Learning.
@@
 By running the policy for $L$ steps, we can obtain the simulated motion sequence $\bm{\hat{x}}^{1:L}$.
-In implementation, we use reinforcement learning to train the motion imitator $\pi$, where a reward is assigned based on how well the simulated motion $\bm{\hat{x}}^{1:L}$ aligns with input motion $\bm{\tilde{x}}^{1:L}$.
+In implementation, we adopt a standard RL algorithm Proximal Policy Optimization (PPO) to train the motion imitator $\pi$, where a reward is assigned based on how well the simulated motion $\bm{\hat{x}}^{1:L}$ aligns with input motion $\bm{\tilde{x}}^{1:L}$.
 In the following, we elaborate on the design of states, rewards, policy and actions.
@@
 Formally, the state is defined as $\bm{s}^{l}=\left[\bm{\tilde{\theta}}^{l+1},\bm{\tilde{p}}^{l+1},\bm{\tilde{\theta}}^{l+1}-\bm{\hat{\theta}}^{l},\bm{\tilde{p}}^{l+1}-\bm{\hat{p}}^{l},\bm{\tilde{v}}^{l+1}-\bm{\hat{v}}^{l},\bm{\tilde{\omega}}^{l}-\bm{\hat{\omega}}^{l}\right]$.
-Here, $\bm{\theta}$, $\bm{p}$, $\bm{v}$ and $\bm{\omega}$ represent the joint rotation, position, joint velocity, and angular velocity, respectively, $\widetilde{\left[\bm{\cdot}\right]}/\widehat{\left[\bm{\cdot}\right]}$ denote the quantities of the input///simulated motions, and $\left[\bm{\cdot}\right]^{l}$ denotes the quantities at the $l^{th}$ timestep.
+Here, $\bm{\theta}$, $\bm{p}$, $\bm{v}$ and $\bm{\omega}$ represent the joint rotation, position, joint velocity, and angular velocity, respectively, $^\widetilde{\left[\bm{\cdot}\right]}$$\widehat{\left[\bm{\cdot}\right]}$ denote the quantities of the input//simulated motions, and $\left[\bm{\cdot}\right]^{l}$ denotes the quantities at the $l^{th}$ timestep.
 Actions.
@@
 The reward function is designed to encourage the simulated motion to match the input motion.
-At each timestep $l$, the reward $r^{l}$ consists of a mimic reward $r^{l}_{\mathrm{m}}$, a energy penalty $r^{l}_{\mathrm{e}}$ and an adversarial reward $r^{l}_{\mathrm{a}}$, formulating as $r^{l}=r^{l}_{\mathrm{m}}+r^{l}_{\mathrm{e}}+r^{l}_{\mathrm{a}}$.
-The mimic reward $r^{l}_{\mathrm{m}}$ measures the difference between the simulated pose and input pose across multiple aspects, including joint rotation $\bm{\theta}$, position $\bm{p}$, joint velocity $\bm{v}$ and angular velocity $\bm{\omega}$, as follows:
+At each timestep $l$, the reward $r^{l}$ consists of a mimic reward $r^{l}_{\mathrm{m}}$ and a energy penalty $r^{l}_{\mathrm{e}}$ and a prior reward $r^{l}_{\mathrm{p}}$, formulating as $r^{l}=r^{l}_{\mathrm{m}}+r^{l}_{\mathrm{e}}+r^{l}_{\mathrm{p}}$.
+The mimic reward $r^{l}_{\mathrm{m}}$ includes joint rotation $\bm{\theta}$, position $\bm{p}$, joint velocity $\bm{v}$ and angular velocity $\bm{\omega}$, as follows:
 $r^{l}_{\mathrm{m}}=w_{\theta}\exp\left[{-\alpha_{\theta}\left|\bm{\tilde{\theta}}^{l}-\bm{\hat{\theta}}^{l}\right|}\right]+w_{p}\exp\left[{-\alpha_{p}\left|\bm{\tilde{p}}^{l}-\bm{\hat{p}}^{l}\right|}\right]$ (2)
 $+w_{v}\exp\left[{-\alpha_{v}\left|\bm{\tilde{v}}^{l}-\bm{\hat{v}}^{l}\right|}\right]+w_{\omega}\exp\left[{-\alpha_{\omega}\left|\bm{\tilde{\omega}}^{l}-\bm{\hat{\omega}}^{l}\right|}\right],$
-[…]
-The energy penalty $r^{l}_{\mathrm{e}}$ regulates the policy by discouraging high-frequency foot jitter, a common issue in policy trained without external forces [5].
-This energy penalty is computed as:
+where $w_{\theta}$, $w_{p}$, $w_{v}$, $w_{\omega}$, are weighting factors, $\alpha_{\theta}$, $\alpha_{p}$, $\alpha_{v}$, $\alpha_{\omega}$ are scaling factors, $^\widetilde{\left[\bm{\cdot}\right]}$$\widehat{\left[\bm{\cdot}\right]}$ denote the quantities of the input//simulated pose, and $\left|\cdot\right|$ is the L1 norm.
+The energy penalty is computed as:
 $r^{l}_{\mathrm{e}}=-0.0005\cdot\left\|\bm{\hat{\nu}}^{l}\bm{\hat{\omega}}^{l}\right\|_{2}^{2},$ (3)
 where $\bm{\hat{\nu}}^{l}$ and $\bm{\hat{\omega}}^{l}$ denotes the joint torque and angular velocity of the simulated pose in the $l^{th}$ timestep, and $\left\|\cdot\right\|_{2}$ is the L2 norm.
-The adversarial reward $r^{l}_{\mathrm{a}}$ encourages the distribution of simulated motions to align that of input motions through a motion discriminator.
-Given the state of current simulated pose, denoted as $\bm{\hat{s}}_{\mathrm{a}}^{l}=[\bm{\hat{\theta}}^{l},\bm{\hat{p}}^{l},\bm{\hat{v}}^{l},\bm{\hat{\omega}}^{l}]$, the adversarial reward is calculated as:
-$r^{l}_{\mathrm{a}}=-\log\left(1-D_{\psi}\left(\bm{\hat{s}}_{\mathrm{a}}^{l}\right)\right).$ (4)
-where $D_{\psi}$ represents the motion discriminator with parameters $\psi$.
-Loss.
-Given the reward $r^{l}$ at each timestep $l$, we train the motion imitator using reinforcement learning (RL).
-The objective is to maximize the excepted reward, effectively enabling the imitator to closely mimic the input motions.
-To achieve this, we adopt a standard RL algorithm Proximal Policy Optimization (PPO) [28] to solve for the optimal imitator policy.
-The objective function is defined as:
-$\mathcal{L}_{\text{PPO}}\left(\phi\right)=\mathbb{E}\left[\min\left(\eta(\phi)r^{l},\operatorname{clip}\left(\eta(\phi),1-\gamma,1+\gamma\right)r^{l}\right)\right].$ (5)
-Here, $\eta(\phi)$ is the probability ratio $\frac{\pi_{\phi}\left(\bm{a}^{l}\mid\bm{s}^{l}\right)}{\pi_{\phi_{\mathrm{old}}}\left(\bm{a}^{l}\mid\bm{s}^{l}\right)}$, where $\phi$ and $\phi_{\mathrm{old}}$ refer to the parameters of the motion imitator after and before the policy update, $\operatorname{clip}$ is an operation that restricts $\eta(\phi)$ within range $\left[1-\gamma,1+\gamma\right]$ with a hyperparameter $\gamma$.
-Motion Discriminator Learning.
-We observe that relying solely on physics simulator often produces motions that feel overly mechanical and unnatural.
-To improve the naturalness of simulated motions, following [22], we introduce a motion discriminator to align the distribution of simulated motions with that of the input motions.
-Specifically, the motion discriminator $D_{\psi}$ learns to distinguish between the state sampled from input pose ($\bm{\tilde{s}}_{\mathrm{a}}^{l}$) and that of simulated pose ($\bm{\hat{s}}_{\mathrm{a}}^{l}$), with the objective function as:
-$\mathcal{L}_{\text{Dis}}\left(\psi\right)$ $=-\mathbb{E}\left[\log\left(D_{\psi}\left(\bm{\tilde{s}}_{\mathrm{a}}^{l}\right)\right)\right]-\mathbb{E}\left[\log\left(1-D_{\psi}\left({\bm{\hat{s}}_{\mathrm{a}}^{l}}\right)\right)\right],$ (6)
-where $\psi$ denotes the parameters of the motion distribution.
+Prior Reward Module.
+Relying solely on physical simulators and imitation learning often results in non-human-like actions, characterized by a mechanical feel and unrealistic rotations.
+Previous works [25, 21] have attempted to address these issues by employing adversarial-based methods trained on high-quality motion data to incorporate human motion knowledge, improving the naturalness of the motion.
+However, this adversarial-based method often suffers from training instability and produces rewards that are not smooth enough, which also prolongs the overall imitation learning training process.
+To address this issue, we propose a Motion VAE Reward model as a prior to constrain the motion generated by the physics simulator, making it more natural.
+As shown in Fig.
+2, to accelerate the training speed of motion imitation learning, we first train a lightweight Motion VAE model using motion data generated by the motion generator.
+This allows the model to effectively learn the distribution of the generated data.
+The objective function for optimizing the Motion VAE model is as follows:
+$\mathcal{L}(\theta,\phi;\mathbf{s})=\mathbb{E}_{q_{\phi}(\mathbf{z}|\mathbf{\Delta s})}\left[\log p_{\theta}(\mathbf{\Delta s}|\mathbf{z})\right]-D_{\text{KL}}\left(q_{\phi}(\mathbf{z}|\mathbf{\Delta s})\parallel p(\mathbf{z})\right)$ (4)
+where $\theta$ and $\phi$ denote the parameters of VAE, $\Delta s$ refer to the difference in state between two consecutive time steps for reference motion.
+$D_{KL}(\cdot,\cdot)$ refers to the KL divergence.
+As shown in Fig. 2, after the Motion VAE model is trained, the pre-trained Motion VAE model is applied to compute rewards as follows:
+[…]
+where $\Delta s$ and $\Delta\overline{s}$ refer to the difference in state between two consecutive time steps for simulated motion and the reconstructed simulated motion, respectively.
+$\left\|\cdot\right\|$ is the L1 norm.
 Initialization and Early Termination.
-During training, we employ reference state initialization [21], where a starting point is randomly selected from a motion clip for imitation.
-We observe that the character frequently falls in the early stages of training.
-To improve training efficiency and accelerate convergence, we adopt an early stopping strategy.
-Specifically, we terminate the episode when the MPJPE (Mean Per Joint Position Error) between the state of input pose and that of simulated pose exceeds $0.5$ meter.
+Similar to [24], a starting point is randomly selected from a motion clip for imitation.
+To improve training efficiency and accelerate convergence, we terminate the episode when the MPJPE (Mean Per Joint Position Error) between the state of input pose and that of simulated pose exceeds $0.5$ meter.
 Hard Negative Mining.
@@
 To address this, we implement a Hard Negative Mining process that identifies motions where the physical simulator fails to imitate as hard samples.
-Specifically, a dynamic weight is assigned to each motion sequence in the input data, doubling whenever imitation fails.
+Specifically, a dynamic weight is assigned to each motion sequence for sampling, doubling whenever imitation fails.
 This process progressively increases the focus on challenging samples, guiding the imitator to effectively learn from difficult examples.
@@
 However, collecting large-scale real 3D motion data is costly.
-To overcome this challenge, we leverage a large volume of synthetic, physically-plausible motion data provided by MPR module, to further enhance the capabilities of the motion generator.
+So, we leverage a large volume of synthetic, physically-plausible motion data provided by MPR module, to further enhance the capabilities of the motion generator.
 As shown in Fig.
@@
 where $\xi$ is the parameters of the motion generator.
+Multi-Round Iterative Optimization.
+In Fig.
+2, we define the complete two-stage process as a single training round.
+To maximize the synergy between the MG and MPR modules, we repeat this process multiple times, using the model weights from each round as initialization for the next, thus creating a multi-round iterative training paradigm.
+We present the results of this iterative training in the Appendix, demonstrating its effectiveness and potential.
 Inference.
@@
 For music-to-dance, we combine Morph with the diffusion-based model EDGE [35] and autoregressive model Bailando [31].
-The Appendix provides additional visualizations and experimental results showcasing Morph’s performance on varying amounts of noisy motion data, different threshold $\tau$ in imitation selection opeartion, and multi-round optimization of the MPR module and motion generator.
+The Appendix provides additional visualizations and experimental results showcasing Morph’s performance on varying amounts of noisy motion data, different and threshold $\tau$ in imitation selection opeartion, and multi-round optimization of the MPR module and motion generator.
 Table 1: Ablation study on Morph-MoMask (combined with MoMask [6] generator) for text-to-motion task on HumanML3D dataset.
-IS: imitation selection operation; Adversarial: using adversarial reward training MPR module; Energy: using energy reward training MPR module; Real Data: using real motion data training MPR module; FT: fine-tuning Motion Generator with physics-refined motions.
-The arrows ($\uparrow$/$\downarrow$ / ↓) indicate that higher/smaller values are better.
+IS: imitation selection operation; Prior: using prior reward training MPR module; Adv: using adversarial reward training MPR module; Energy: using energy reward training MPR module; Real Data: using real motion data training MPR module; FT: fine-tuning Motion Generator with physics-refined motions.
+The arrows ($\uparrow/\downarrow$) indicate that higher/smaller values are better.
 Model Number Methods Common Generation Metrics Physical Plausibility Metrics
-IS Adversarial Energy Real Data FT RTOP-3 $\uparrow$ FID $\downarrow$ PFC $\downarrow$ Penetrate $\downarrow$ Float $\downarrow$ Skate $\downarrow$ IFR $\downarrow$
+IS Prior Adv Energy Real Data FT RTOP-3 $\uparrow$ FID $\downarrow$ PFC $\downarrow$ Penetrate $\downarrow$ Float $\downarrow$ Skate $\downarrow$ IFR $\downarrow$
 $\mathrm{A}$ Baseline (only motion generator MoMask [6]) 0.807 0.045 1.058 23.152 10.660 5.262 -
-$\mathrm{B}$ 0.792 0.194 0.852 0.000 2.272 0.020 0.0155
-$\mathrm{C}$ 0.782 0.276 0.735 0.000 2.554 0.032 0.0362
-$\mathrm{D}$ 0.790 0.165 0.715 0.000 2.376 0.026 0.0272
-$\mathrm{E}$ 0.801 0.077 0.672 0.000 2.271 0.011 0.0155
-$\mathrm{F}$ 0.785 0.183 0.749 0.000 2.451 0.017 0.0338
-$\mathrm{G}$ 0.816 0.043 0.651 0.000 2.146 0.010 0.0151
+$\mathrm{B}$ 0.800 0.172 0.821 0.000 2.233 0.016 0.0152
+$\mathrm{C}$ 0.792 0.194 0.852 0.000 2.272 0.020 0.0155
+$\mathrm{D}$ 0.782 0.276 0.735 0.000 2.554 0.032 0.0362
+$\mathrm{E}$ 0.795 0.163 0.711 0.000 2.362 0.024 0.0270
+$\mathrm{F}$ 0.802 0.074 0.669 0.000 2.268 0.011 0.0153
+$\mathrm{G}$ 0.792 0.179 0.734 0.000 2.420 0.015 0.0327
+$\mathrm{H}$ 0.816 0.041 0.647 0.000 2.141 0.010 0.0149
 4.1 Datasets and Evaluation Metrics
@@
 We evaluate motion generation from two perspectives: common generation metrics and physical plausibility metrics.
-We first present the common generation metrics.
-For text-to-motion generation, following [8, 41], we use three standard metrics: Frechet Inception Distance (FID) that measures the distance between the generated and ground-truth motion distributions; R-Precition (including RTOP-1, RTOP2, and RTOP-3) that computes the retrieval accuracy of the generated motions with respect to the input text; Diversity (Div) that measures the diversity of generated motions.
-For music-to-dance generation, following [16, 38, 29], we use five standard metrics: FIDk and Divk that calculate Frechet Inception Distance and Diversity on kinetic features (denoted as ‘$k$’); FIDg and Divg that calculate Frechet Inception Distance and Diversity on geometric features (denoted as ‘$g$’); Beat Align Score (BAS) that measures the alignment between input music and generated dances.
+First, for text-to-motion generation, we follow [7, 43] and use three standard metrics: Frechet Inception Distance (FID), which measures the distance between generated and ground-truth motion distributions; R-Precision (RTOP-1, RTOP-2, RTOP-3), which computes retrieval accuracy of generated motions based on input text; and Diversity (Div), which measures the diversity of generated motions.
+For music-to-dance generation, following [15, 40, 31], we use five metrics: FIDk and Divk for kinetic features (denoted ‘$k$’); FIDg and Divg for geometric features (denoted ‘$g$’); and Beat Align Score (BAS), which measures the alignment between input music and generated dances.
 Next, we present the physical plausibility metrics.
@@
 The NVIDIA’s Isaac Gym is used as the physics simulator.
-Rather than using real motion data, we train Morph on synthetic motion data produced by a pre-trained model, increasing the motion data volume to three times the original.
-In the training phase of MPR module (Stage 1), we use the Adam optimizer with a batch size of $64$ and a learning rate of $4\times 10^{-5}$.
+During the MPR module’s training phase (Stage 1), we employ the Adam optimizer, with a batch size of 64 and a learning rate of $4\times 10^{-5}$.
 During the finetuning of motion generator (Stage 2), we follow the original training setup, only adjusting the learning rate to $1\times 10^{-5}$.
@@
 2) to $0.5$, $0.3$, $0.1$ and $0.1$, $\alpha_{\theta}$, $\alpha_{q}$, $\alpha_{v}$ and $\alpha_{w}$ (Eq.
-2) to $100$, $10$, $0.1$ and $0.1$, $\gamma$ (Eq.
-5) is set to $0.1$ and the imitation selection threshold $\tau$ is set to $0.5$.
+2) to $100$, $10$, $0.1$ and $0.1$ and the imitation selection threshold $\tau$ is set to $0.5$.
 All our experiments are conducted on 8 Tesla V100 GPUs.
@@
 Therefore, we annotate each text description in HumanML3D [8] based on its semantic content.
-Text Descriptions involving interaction-based motions are labeled as ‘00’, while other descriptions are labeled as ‘$1$’.
+Text Descriptions involving interaction-based motions are labeled as ‘0’, while other descriptions are labeled as ‘$1$’.
 Text labeled as ‘$1$’ will be used to train the MPR module in Stage 1.
@@
 […]
-Morph-MDM 0.467 0.752 0.485 9.863 0.707 0.000 2.261 0.018 0.0154
+Morph-MDM 0.471 0.757 0.482 9.865 0.697 0.000 2.258 0.016 0.0153
 MotionDiffuse(MD) [42] - - 0.491 0.782 0.630 9.410 0.533 16.670 5.698 3.419 -
 […]
-Morph-MD 0.494 0.786 0.571 9.492 0.416 0.000 0.856 0.031 0.0166
+Morph-MD 0.494 0.788 0.568 9.493 0.414 0.000 0.853 0.029 0.0163
 T2M-GPT [42] - - 0.491 0.775 0.116 9.761 0.998 72.250 8.918 7.801 -
 […]
-Morph-T2M-GPT 0.497 0.781 0.109 9.768 0.751 0.000 2.703 0.042 0.0197
+Morph-T2M-GPT 0.498 0.784 0.105 9.771 0.748 0.000 2.700 0.039 0.0195
 MoMask [6] - - 0.521 0.807 0.045 9.641 1.058 23.152 10.660 5.262 -
 […]
-Morph-MoMask 0.525 0.816 0.043 9.687 0.651 0.000 2.146 0.010 0.0151
+Morph-MoMask 0.525 0.816 0.041 9.689 0.645 0.000 2.141 0.010 0.0149
 Table 3: Comparison results on common generation metrics for text-to-motion on HumanML3D dataset.
-Methods RTOP-1 $\uparrow$ RTOP-2 $\uparrow$ RTOP-3 $\uparrow$ FID $\downarrow$ MM-Dist $\downarrow$ Diversity $\uparrow$ MModality $\uparrow$
+Methods RTOP-1 $\uparrow$ RTOP-2 $\uparrow$ RTOP-3 $\uparrow$ FID $\downarrow$ $M$-Dist $\downarrow$ Diversity $\uparrow$ MModality $\uparrow$
 MDM [33] 0.455 0.645 0.749 0.489 3.330 9.920 2.290
@@
 BAMM [26] 0.525 0.720 0.814 0.055 2.919 9.717 1.687
-Morph-MDM 0.467 0.658 0.752 0.485 3.106 9.863 2.177
-Morph-MD 0.494 0.687 0.786 0.571 3.001 9.492 1.584
-Morph-T2M-GPT 0.497 0.684 0.781 0.109 2.997 9.768 1.903
-Morph-MoMask 0.525 0.722 0.816 0.043 2.923 9.687 1.652
+Morph-MDM 0.471 0.664 0.757 0.482 3.103 9.865 2.179
+Morph-MD 0.494 0.691 0.788 0.568 3.000 9.494 1.582
+Morph-T2M-GPT 0.498 0.690 0.784 0.105 2.994 9.771 1.910
+Morph-MoMask 0.525 0.726 0.816 0.041 2.920 9.689 1.655
 Table 4: Comparison results for music-to-dance on AIST++ dataset. $\dagger$ denotes Morph without fine-tuning motion generator.
@@
 […]
-Morph-EDGE 39.29 19.88 4.91 5.23 0.2327 0.287 0.000 3.524 0.010 0.0085
+Morph-EDGE 39.24 19.82 4.93 5.36 0.2329 0.295 0.000 3.519 0.009 0.0082
 […]
-Morph-Bailando 26.32 9.43 7.86 6.55 0.2411 0.045 0.000 2.061 0.023 0.0059
-The generated motion sequences may exhibit issues such as body leaning, floating and ground penetration.
-When imported into the engine, these issues can cause instability in the robot, leading to falls, bouncing off the ground, or dropping from mid-air.
-To address this, we apply a preprocessing step to the motion sequences.
-Specifically, we first compute the body’s tilt angle, defined as the angle between the projection of the center of mass onto the ground and the line connecting both feet, and apply this angle to the pelvis for the entire sequence.
-To correct for floating and penetration, we determine the lowest point height and adjust the entire sequence by this offset.
-The resulting sequence is then used for training and inference.
-More details can be found in the Appendix.
+Morph-Bailando 26.25 9.38 7.88 6.62 0.2412 0.043 0.000 2.057 0.021 0.0055
 4.3 Ablation Studies
@@
 We investigate the effectiveness of using only synthetic motion data from two perspectives: its overall effectiveness and its performance relative to training with real motion data.
-First, we compare the baseline model $\mathrm{A}$ (only the MoMask generator) and model $\mathrm{E}$ which combines MoMask with the trained MPR module.
-As shown in Tab.
-1, model $\mathrm{E}$ outperforms model $\mathrm{A}$ on all physical metrics ($0.651/1.058$ on PFC, $0.0/23.152$ on Penetrate, $2.146/10.66$ on Float, $0.01/5.262$ on Skate), while maintaining competitive performance on generation metrics.
+First, we compare the baseline model $\mathrm{A}$ (only the MoMask generator) and model $\mathrm{F}$ which combines MoMask with the trained MPR module.
+In Tab.
+1, model $\mathrm{F}$ outperforms model $\mathrm{A}$ on all physical metrics ($0.669/1.058$ on PFC, $0.0/23.152$ on Penetrate, $2.268/10.66$ on Float, $0.011/5.262$ on Skate), while maintaining competitive performance on generation metrics.
 These results confirm the feasibility of training the MPR module solely with synthetic data.
-Next, we compare model $\mathrm{E}$ with model $\mathrm{F}$, which uses real motion data to train the MPR module.
-As shown in Tab.
-1, model $\mathrm{F}$ achieves inferior performance compared to model $\mathrm{E}$.
+Next, we compare model $\mathrm{F}$ with model $\mathrm{G}$, which uses real motion data to train the MPR module.
+In Tab.
+1, model $\mathrm{G}$ achieves inferior performance compared to model $\mathrm{F}$.
 This can be attributed to the significant domain gap between the training data (real motions) and test data (generated motions).
 This gap hinders the MPR module trained on real data from effectively adapting to the generated motion data, leading to performance decline.
-Effectiveness of adversarial reward and energy reward.
-When training the MPR module, we add an adversarial reward and energy reward in addition to the common imitation task reward.
+Effectiveness of prior reward and energy reward.
+When training the MPR module, we add an prior reward and energy reward alongside the common imitation task reward.
 In Tab.
-1, comparing model $\mathrm{E}$ with model $\mathrm{C}$, which omits the adversarial reward, we observe a performance decline in both generation and physical metrics on model $\mathrm{C}$.
-The significant drop in FID metric indicates that, without the adversarial reward, the simulated motions lack the distribution constraints of the input motions, leading to a deviation from the intended motion distribution.
-Next, we compare model $\mathrm{E}$ with $\mathrm{D}$, which omits the energy reward.
+1, comparing model $\mathrm{F}$ with model $\mathrm{D}$, which omits the prior reward, we observe a performance decline in both generation and physical metrics on model $\mathrm{D}$.
+The significant drop in FID metric indicates that, without the prior reward, the simulated motions lack the distribution constraints of the input motions, leading to a deviation from the intended motion distribution.
+The $\mathrm{B}$, which leverages prior rewards, significantly outperforms model $\mathrm{C}$, which uses adversarial rewards, across various generation and physics-based metrics.
+This highlights the superior effectiveness of the proposed prior reward over the adversarial reward.
+We compare model $\mathrm{F}$ with $\mathrm{E}$, which omits the energy reward.
 As shown in Tab.
-1, model $\mathrm{E}$ outperforms model $\mathrm{D}$ in generation metrics.
+1, model $\mathrm{F}$ outperforms model $\mathrm{E}$ in generation metrics.
 We argue that introducing energy reward helps suppress high-frequency jitter during the humanoid robot control process, resulting in more natural motions.
@@
 In Tab.
-1, we compare model $\mathrm{E}$ with $\mathrm{B}$, which omits the imitation selection operation.
-As shown, model $\mathrm{E}$ outperforms $\mathrm{B}$ in both generation and physic metrics, validating the effectiveness of the imitation selection strategy.
+1, we compare model $\mathrm{F}$ with $\mathrm{B}$, which omits the imitation selection operation.
+As shown, model $\mathrm{F}$ outperforms $\mathrm{B}$ in both generation and physic metrics, validating the effectiveness of the imitation selection strategy.
 Effectiveness of finetuning motion generator using simulated motions.
-Furthermore, we compare model $\mathrm{E}$ and model $\mathrm{G}$, which finetunes the motion generator using the physically refined motions produced by the MPR module.
+Furthermore, we compare model $\mathrm{F}$ and model $\mathrm{H}$, which finetunes the motion generator using the physically refined motions produced by the MPR module.
 As shown in Tab.
-1, model $\mathrm{G}$ achieves further gains in both generation and physic metrics, with RTOP-3 increasing by $1.5\%$ and FID improving by $0.034$.
+1, model $\mathrm{H}$ achieves further gains in both generation and physic metrics, with RTOP-3 increasing by $1.4\%$ and FID improving by $0.033$.
 These results suggest that the proposed MPR module can, in turn, enhance the performance of the motion generator.
@@
 2, Morph significantly enhances physical fidelity while maintaining competitive generation metrics across different motion generation models.
-For example, after integrating Morph with MDM (Morph-MDM), RTOP-1 metric improved by $0.012$.
-In terms of physical plausibility metrics, penetration dropped to zero, float decreased from $17.502$ to $2.261$, and skate reduced from $3.540$ to $0.018$.
+For example, after integrating Morph with MDM (Morph-MDM), RTOP-1 metric improved by $0.016$.
+In terms of physical plausibility metrics, penetration dropped to zero, float decreased from $17.502$ to $2.258$, and skate reduced from $3.540$ to $0.016$.
 Similar performance improvements are observed with other motion generators.
@@
 4, Morph significantly improves physical metrics compared to existing methods.
-In terms of generation metrics, our Morph-EDGE model achieves the best results in multiple metrics, including FIDg, Divg, and BAS, while also demonstrating competitive performance in FIDk metric.
+For generation metrics, our Morph-EDGE model achieves the best results in multiple metrics, including FIDg, Divg, and BAS, while also demonstrating competitive performance in FIDk.
 These results validate the superiority of our framework in music-to-dance generation.
 5 Conclusion
-In this paper, we present Morph, a model-agnostic physical optimization framework, designed to enhance physical plausibility in motion generation without relying on costly real-world motion data.
-To accomplish this, we first leverage a pretrained motion generator to synthesize large-scale noisy motion data.
-We then introduce a Motion Physics Refinement module, which utilizes these synthesized data to train a motion imitator that enforces physical constraints.
-The physically refined motions can, in turn, be used to finetune the motion generator, further enhancing its capabilities.
-Our framework is compatible with various motion generation models across both text-to-motion and music-to-dance generation tasks.
-Extensive experimental results demonstrate that Morph substantially improves physical plausibility while achieving competitive generation quality.
+In this paper, we present Morph, a model-agnostic physical optimization framework designed to enhance physical plausibility in motion generation without relying on costly real-world data.
+We begin by using a pretrained motion generator (MG) to synthesize large-scale noisy motion data.
+Then, we introduce a Motion Physics Refinement (MPR) module that leverages this synthesized data to train a motion imitator enforcing physical constraints.
+These physically refined motions are used to fine-tune the motion generator, further enhancing its capabilities.
+To stabilize the physics optimization process, we introduce a prior reward module.
+This collaborative training paradigm enables mutual enhancement between MG and MPR module, improving practicality and robustness.
+Our framework is compatible with various motion generation models for both text-to-motion and music-to-dance tasks.
+Extensive experiments show that Morph substantially improves physical plausibility while achieving competitive generation quality.
 Appendix
@@
 C), and effect of the number of training rounds for Morph (Sec.
-D).
+E).
 Finally, we provide additional qualitative comparisons for text-to-motion and music-to-dance tasks (Sec.
-E).
+H).
 Appendix A Details for Data Preprocess
@@
 $\tau$=0.0 0.521 0.807 0.045 9.641 1.058 23.152 10.660 5.262 -
-$\tau$=0.1 0.520 0.806 0.051 9.633 0.879 3.567 4.783 2.132 0.1287
-$\tau$=0.2 0.518 0.806 0.058 9.629 0.774 0.842 4.020 1.061 0.0543
-$\tau$=0.3 0.515 0.804 0.071 9.578 0.760 0.058 3.202 0.533 0.0262
-$\tau$=0.4 0.514 0.803 0.073 9.582 0.728 0.003 2.998 0.217 0.0162
-$\tau$=0.5 0.511 0.801 0.077 9.579 0.672 0.000 2.271 0.011 0.0155
-$\tau$=0.6 0.508 0.799 0.080 9.575 0.686 0.000 2.268 0.011 0.0147
-$\tau$=0.7 0.506 0.799 0.081 9.543 0.663 0.000 2.102 0.008 0.0131
-$\tau$=0.8 0.500 0.795 0.081 9.522 0.647 0.000 2.028 0.006 0.0126
-$\tau$=0.9 0.496 0.793 0.087 9.412 0.638 0.000 1.989 0.005 0.0120
-$\tau$=1.0 0.488 0.790 0.088 9.260 0.625 0.000 1.984 0.005 0.0113
+$\tau$=0.1 0.520 0.806 0.048 9.636 0.877 3.564 4.779 2.128 0.1281
+$\tau$=0.2 0.519 0.805 0.056 9.625 0.771 0.838 4.015 1.057 0.0540
+$\tau$=0.3 0.518 0.805 0.067 9.583 0.757 0.054 3.200 0.529 0.0258
+$\tau$=0.4 0.517 0.803 0.071 9.584 0.722 0.002 2.991 0.211 0.0158
+$\tau$=0.5 0.516 0.802 0.074 9.578 0.669 0.000 2.268 0.011 0.0153
+$\tau$=0.6 0.512 0.801 0.079 9.576 0.664 0.000 2.263 0.010 0.0144
+$\tau$=0.7 0.510 0.799 0.080 9.543 0.660 0.000 2.093 0.006 0.0128
+$\tau$=0.8 0.506 0.797 0.081 9.520 0.645 0.000 2.022 0.005 0.0124
+$\tau$=0.9 0.504 0.795 0.085 9.408 0.634 0.000 1.985 0.004 0.0117
+$\tau$=1.0 0.497 0.793 0.084 9.255 0.623 0.000 1.982 0.003 0.0111
 […]
@@
 RTOP-1 $\uparrow$ RTOP-3 $\uparrow$ FID $\downarrow$ Diversity $\uparrow$ PFC $\downarrow$ Penetrate $\downarrow$ Float $\downarrow$ Skate $\downarrow$ IFR $\downarrow$
-One-Round w/o FT 0.511 0.801 0.077 9.579 0.672 0.000 2.271 0.011 0.0155
-One-Round 0.525 0.816 0.043 9.687 0.651 0.000 2.146 0.010 0.0151
-Two-Round w/o FT 0.525 0.818 0.043 9.689 0.635 0.000 2.132 0.010 0.0137
-Two-Round 0.527 0.821 0.044 9.692 0.631 0.000 2.117 0.008 0.0134
+One-Round w/o FT 0.516 0.802 0.074 9.578 0.669 0.000 2.268 0.011 0.0153
+One-Round 0.525 0.816 0.041 9.689 0.645 0.000 2.141 0.010 0.0149
+Two-Round w/o FT 0.526 0.817 0.041 9.692 0.632 0.000 2.129 0.010 0.0134
+Two-Round 0.527 0.818 0.038 9.697 0.625 0.000 2.108 0.007 0.0129
+Three-Round w/o FT 0.526 0.821 0.040 9.701 0.621 0.000 2.121 0.009 0.0131
+Three-Round 0.528 0.823 0.034 9.715 0.618 0.000 2.100 0.006 0.0122
 Appendix B Effect of $\tau$ in Imitation Selection on Morph
@@
 […]
+Figure 5: T-sne of motion and text distribution between MG and Morph.
+Table 8: The win rate of Morph over baselines.
+Module Semantic Consistency Realism Physical Plausibility Fluency
+MDM-Morph vs. MDM 84.8% 80.1% 96.6% 85.0%
+T2M-GPT-Morph vs. T2M-GPT 87.1% 79.6% 94.4% 81.2%
+MoMask-Morph vs. MoMask 90.4% 88.3% 97.5% 80.9%
 Appendix D Effect of Multi-Round Optimization of the MPR module and MG on Morph
@@
   • Two-Round: Fine-tuning the Motion Generator using the results from Two-Round w/o FT.
+  • Three-Round w/o FT: Training the MPR module again using the motion data generated by the fine-tuned MG from the second round.
+  • Three-Round: Fine-tuning the Motion Generator using the results from Three-Round w/o FT.
 As shown in Tab.
@@
 In the second round, the fine-tuned MG from the first round is used to generate training data for the MPR module (initialized with first-round weights).
-We observed improvements in Two-Round w/o FT compared to One-Round, with PFC increasing by 0.016, Float by 0.014, and IFR decreasing, indicating enhanced motion imitation by the MPR module.
+We observed improvements in Two-Round w/o FT compared to One-Round, with PFC increasing by 0.013, Float by 0.012, and IFR decreasing, indicating enhanced motion imitation by the MPR module.
 After fine-tuning the MG once again, Two-Round shows improvements in the RTOP-1 and RTOP-3 metrics.
+The model’s generation and physical performance reached their best in Three-Round.
 These results clearly demonstrate that the MG and MPR modules can mutually enhance each other.
 Moreover, alternating training between the MG and MPR modules across multiple rounds can further improve the performance of Morph.
-Appendix E More Qualitative Results
+Table 9: Cross-Task generalization results on Music2Dance and Text2Motion
+Task $FID/FID_{k}$ $\downarrow$ $FID_{g}$ $\downarrow$ $RTOP3/Div_{k}$ $\uparrow$ PFCPFC $\downarrow$ PenetratePenetrate $\downarrow$ FloatFloat $\downarrow$
+Cross-task generalization evaluation
+Text-to-Motion (MPR trained on T2M) 0.074 0.802 0.669 0.000 2.268
+Text-to-Motion (MPR trained on M2D) 0.116 0.795 0.881 0.000 2.436
+Music-to-Dance (MPR trained on M2D) 35.48 7.70 12.03 0.044 0.000 2.076
+Music-to-Dance (MPR trained on T2M) 37.66 7.38 14.02 0.057 0.000 2.193
+Action-to-motion tests
+Action_Label-to-Motion (MDM-action) 0.497 0.396 0.544 15.770 7.467
+Action_Label-to-Motion (MDM-action with Morph) 0.424 0.416 0.509 0.000 2.115
+GAN-based Transformer tests
+Text-to-Motion (GAN-based Transformer) 0.628 - 0.736 0.933 47.612 21.008
+Text-to-Motion (GAN-based Transformer with Morph) 0.606 - 0.755 0.742 0.000 2.637
+Long-duration dance samples tests
+Music-to-Dance (30s long-term dance, Lodge) 45.56 34.29 6.75 0.114 46.772 29.857
+Music-to-Dance (30s long-term dance, Lodge-Morph) 43.96 32.88 6.90 0.083 0.000 3.163
+PHC-based baseline
+Text-to-Motion (MoMask+PHC) 0.183 - 0.785 0.749 0.000 2.451
+Text-to-Motion (MoMask+Morph) 0.041 - 0.816 0.647 0.000 2.141
+Appendix E Semantic Alignment Analysis
+As shown in Fig.
+5, Morph significantly outperforms MG on alignment metrics.
+Globally, Morph demonstrates superior realism by closely matching the statistical distribution of real motions—exhibiting similar clustering patterns and range of variation (Left).
+Crucially, it also achieves stronger semantic alignment with input text features, forming tighter clusters around corresponding text embeddings to better capture intended meanings (Middle).
+Locally, Morph provides enhanced semantic matching at the segment level, ensuring fine-grained motion elements correspond more accurately to detailed text semantics throughout the sequence (Right).
+In summary, Morph can generate semantically faithful, realistic motions compared to the MG baseline.
+Appendix F Cross-Task Generalization Ability
+To evaluate the cross-task generalization of the MPR module, we conducted cross-validation by testing its performance across two distinct tasks: text-to-motion (using the MoMask dataset) and music-to-dance (using Bailando).
+These tasks differ significantly in input modalities—one driven by linguistic descriptions, the other by rhythmic audio-and in motion characteristics, from daily actions to stylized dance moves.
+As shown in Tab.
+9, the MPR module retains strong performance even when trained without task-specific synthetic data: it not only preserves motion quality (e.g., smooth transitions and natural postures) but also maintains physical plausibility (avoiding joint distortions or gravity-defying movements).
+This confirms its ability to generalize beyond specific task boundaries.
+Appendix G User Study
+In the user study, we evaluated win rates across four critical dimensions—semantic alignment (matching text descriptions), authenticity (resembling real motions), physical plausibility (avoiding unnatural joint movements), and fluency (smooth temporal transitions).
+Morph decisively outperformed baseline methods here: as shown in Tab.
+8, it achieved significantly higher win rates of 87.4%, 85.2%, 96.2%, and 82.4% respectively.
+Such consistent leads across all key metrics confirm its comprehensive advantages in motion generation quality.
+Appendix H More Qualitative Results
 Fig. 6 and Fig. 7 provide the additional qualitative results for the text-to-motion and music-to-dance generation tasks using Morph.
 As shown in Fig.
-5, in the text-to-motion generation task, floating and penetration are common artifacts in motion generation, often resulting from inaccuracies in the estimation of translation.
+6, in the text-to-motion generation task, floating and penetration are common artifacts in motion generation, often resulting from inaccuracies in the estimation of translation.
 However, Morph effectively addresses these issues, successfully mimicking the input motion and demonstrating a significant improvement in mitigating these artifacts.
@@
 As shown in Fig.
-6, in the music-to-dance generation task, floating and penetration are the most prominent issues.
+7, in the music-to-dance generation task, floating and penetration are the most prominent issues.
 Due to the faster frequency of dance movements, these artifacts occur more frequently.